OpenStudy (anonymous):

What is the value of z in the equation 2(4z − 6 − 6) = 170 − 46? 12.5 15.5 16.5 18.5

OpenStudy (owlcoffee):

well, all you have to do is simplify and solve for "z"

OpenStudy (anonymous):

ik that but i did and i was told i got the wrong asnwer

OpenStudy (owlcoffee):

what did you get?

OpenStudy (anonymous):

12.5

OpenStudy (anonymous):

is that right ?

OpenStudy (owlcoffee):

So, all you have to do is simplify: \[2(4z-6-6)=170-46\] \[8z-12-12=124\] \[8z-24=124\] \[z-3=15,5\] \[z=18,5\]

OpenStudy (anonymous):

ok well that makes more sense thank you !